IAN–CNR

via Abbiategrasso 209 I-27100 Pavia ITALY

Pavia, February 22, 1999

Prof. G.A. Watson Editor-In-Chief The IMA Journal of Numerical Analysis Department of Mathematics Dundee DD1 4HN, Scotland Great Britain

Dear Prof. Watson

Please find here enclosed four copies of the paper “Wavelet Stabilization and Preconditioning for Domain Decomposition”, by A. Kunoth and myself, which we would like to submit for publication on the IMA Journal of Numerical Analysis.

Best regards,

Silvia Bertoluzza Senior Researcher

e-mail: aivlis@dragon.ian.pv.cnr.it